Direct Method of Support Department Cost Allocation Valron Company has two support departments, Human Resources and...

  1. Direct Method of Support Department Cost Allocation

    Valron Company has two support departments, Human Resources and General Factory, and two producing departments, Fabricating and Assembly.


    Support Departments
    Producing Departments
    Human
    Resources
    General
    Factory
    Fabricating Assembly
    Direct costs $150,000 $330,000 $114,200 $94,000
    Normal activity:
    Number of employees 60 80 170
    Square footage 1,000 5,700 13,300

    The costs of the Human Resources Department are allocated on the basis of number of employees, and the costs of General Factory are allocated on the basis of square footage. Valron Company uses the direct method of support department cost allocation.

    Required:

    1. Calculate the allocation ratios for the four departments using the direct method. If an amount is zero, enter "0". Round your answer to the nearest cent.

    Proportion of Driver Used by
    Human Resources General Factory Fabricating Assembly
    Human Resources
    General Factory

    2. Using the direct method, allocate the costs of the Human Resources and General Factory departments to the Fabricating and Assembly departments. If an amount is zero, enter "0".

    Support Departments Producing Departments
    Human Resources General Factory Fabricating Assembly
    Direct costs $ $ $ $
    Allocate:
      Human Resources            
      General Factory            
    Total after allocation $ $ $ $
Answer #1

Answers

  • Requirement a

Proportion of Driver Used by

Human Resources

General Factory

Fabricating

Assembly

Human Resources

                        -  

                                  -  

                         0.3200

                       0.6800

General Factory

                        -  

                                  -  

                                 0.3000

                                0.7000

--Working

Proportion of Driver Used by

Human Resources

General Factory

Fabricating

Assembly

Human Resources

0

0

=80/(80+170)

=170/(80+170)

General Factory

0

0

=5700/(5700+13300)

=13300/(5700+13300)

  • Requirement [b]

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

$150,000

$330,000

$114,200

$94,000

Allocate:

Human Resources

($150,000)

$48,000

$102,000

General Factory

$0

($330,000)

$99,000

$231,000

Total after allocation

$0

$0

$261,200

$427,000

--Working

Support Departments

Producing Departments

Human Resources

General Factory

Fabricating

Assembly

Direct costs

150000

330000

114200

94000

Allocate:

Human Resources

-150000

=150000*0.32

=150000*0.68

General Factory

-330000

=330000*0.3

=330000*0.7